ganeshie8 (ganeshie8):

whats the difference between the function compositions \((f\circ g)\circ h\) and \(f \circ (g\circ h)\)

OpenStudy (welshfella):

is there a difference?

ganeshie8 (ganeshie8):

It turns out the compositions are associative, but what exactly is the difference between them ?

OpenStudy (zzr0ck3r):

Hint: Function composition is associative

OpenStudy (ikram002p):

there is no difference only be careful about range and domain

ganeshie8 (ganeshie8):

i mean, in terms of execution order... what difference that parenthesis ordering make ? idk im having a hard time moving forward with my review of abstract algebra because of this..

OpenStudy (ikram002p):

order might affect on that only.

ganeshie8 (ganeshie8):

Hmm okay, domain and codomain are valid in both cases ``` h : X -> Y g : Y -> Z f : Z -> W ```

OpenStudy (freckles):

hey so you do you mean like what would \[f \circ (g \circ h) \text{ verus what would it look like to find } (f \circ g) \circ h\] \[f=x^2 \\ g=x-1 \\ h= \sqrt{x} \\ f \circ g =(x-1)^2 \\ g \circ h=\sqrt{x}-1 \\ ( f \circ g) \circ h=(\sqrt{x}-1)^2 \\ f \circ (g \circ h)=(\sqrt{x}-1)^2\]?

ganeshie8 (ganeshie8):

I see compositions don't commute in general : \(f\circ g \ne g\circ f\) But I am studying associativity as of now. for example, we knw that addition over real numbers is associative : ``` (a + b) + c = a + (b + c) ``` here we agree that "a+b" is executed first on left hand side. is there a similar thing for compositions ?

OpenStudy (ikram002p):

take this example h : {b,1,2,3}->{1 } g : {1,2,3}->{4,5,a,c} f : {.4,1,2,3}->{a,b,c}

OpenStudy (freckles):

yes you can find f composed with g first then plug in h or you can find g composed with h then plug that result into f

ganeshie8 (ganeshie8):

how does that work ? say \(f(x) = x\) \(g(x) = 2x\) \(h(x)=3x\)

ganeshie8 (ganeshie8):

\((f\circ g)\circ h(2) = ?\) what does it mean to compose \(f\) and \(g\) first, here ?

OpenStudy (freckles):

\[(f \circ g)(x)=2x \\ (g \circ h)(x)=2(3x) \\ \text{ now we will replace } x \text{ in first one with } h(x) \\ (f \circ g)(h(x))=2(3x) \\ \text{ now we will replace the } x \text{ in } f(x)=x \text{ with } (g \circ h)(x) \\ \text{ so we have } \\ f((g \circ h)(x))=2(3x)\]

ganeshie8 (ganeshie8):

I think I get it! thanks a lot for breaking it down !

OpenStudy (freckles):

I kind of like to leave the x off sometimes

OpenStudy (freckles):

\[(f \circ g) \circ h \\ \text{ where } f=x^2 \\ g=x-1 \\ h=\sqrt{x} \\ f \circ g =(x-1)^2 \\ \text{ then the input is } h \\ (f \circ g) (h)=(\sqrt{x}-1)^2\] it seems less confusing to me for some reason without writing the x

ganeshie8 (ganeshie8):

I see... guess I never played with these much, its kinda fun reviewing/learning these now :) Was going through this proof earlier when I got stuck on compositions : It can be shown that every square matrix represents a composition, therefore the matrix multiplication is also associative!

OpenStudy (zzr0ck3r):

\(f\) is the function \(f(x) = 3x+2\) is the rule \(f(x)\) does not really make sense, but we abuse the hell out of the notation :)

ganeshie8 (ganeshie8):

f : x -> 3x+2 is more proper i think, hope ill be able to appreciate these fine details after finishing abstract algebra xD

OpenStudy (ikram002p):

ur doing abstract now ?

OpenStudy (freckles):

what?!? you really are doing abstract algebra?

OpenStudy (zzr0ck3r):

\(f:\mathbb{R}\rightarrow \mathbb{R}, f(x) = 3x+2\) is a function

OpenStudy (freckles):

I thought you already graduated and stuff. You know everything. :p

OpenStudy (zzr0ck3r):

\(f:\underset{x\mapsto 3x+2}{\mathbb{R}\rightarrow \mathbb{R}}\)

OpenStudy (zzr0ck3r):

I really like that last one, but it takes so long in latex

OpenStudy (ikram002p):

hmmm yeah using this notation is better

OpenStudy (zzr0ck3r):

I would not call this abstract algebra. Everything in math is a set. Almost everything is a function.

ganeshie8 (ganeshie8):

lol i suck at math, i am an engineer, what do you think haha! I think abstract algebra is more like pure mathematics , never had to touch this before.

OpenStudy (ikram002p):

nah some times u are perfect in math gane.

ganeshie8 (ganeshie8):

Ahh nice, is it correct to interpret that as : \(f\) is a relation between the sets \(\mathbb R\) and \(\mathbb{R}\) such that \(x\) maps to \(3x+2\)

OpenStudy (zzr0ck3r):

I would say that \(f\) is the name of the function, the name of the rule that gives the relation is \(f(x) = 2x\).

OpenStudy (zzr0ck3r):

so we could say \(f= \{(1,2),(2,4),(4,5)\}\) but \(f(x)= \{(1,2),(2,4),(4,5)\}\) makes no sense.

OpenStudy (zzr0ck3r):

and in this situation we could say \(f(\{1,2,4\})=\{2,4,5\}\).

OpenStudy (zzr0ck3r):

some people would want \(f[\{1,2,4\}]=\{2,4,5\}\)

OpenStudy (ikram002p):

i thought this is also correcto \(f (\{(1,2),(2,4),(4,5)\} )\)

OpenStudy (anonymous):

\((f\circ g)(x)=f(g(x))\\((f\circ g)\circ h)(x)=(f\circ g)(h(x))=f(g(h(x)))\\(f\circ(g\circ h))(x)=f((g\circ h)(x))=f(g(h(x)))\)

OpenStudy (anonymous):

this is the 'difference' in order, @ganeshie8